How many atoms of Oxygen has are there in 120.0L?


1
Expert's answer
2022-03-16T23:10:03-0400

120L= 120kg= 120000

I mole of oxygen = 16g

Moles= 120000/16= 7500moles

Avagadro constant= 6.022× 10^23

Atoms= 4.5× 10^27